OpenStudy (dtan5457):

Solve for sin a/2 when cos a=4/5 , o

OpenStudy (dtan5457):

just to check, i got \[\frac{ 3\sqrt{10} }{ 10 }\] correct?

OpenStudy (dtan5457):

@Hoslos @Michele_Laino @mathstudent55

OpenStudy (dtan5457):

oh crap why is this in chemistry...

OpenStudy (anonymous):

Right, using the calculator to first find a, you do :\[a=\cos^{-1} 4/5=36.9\] Then do\[\sin (36.9/2)=0.316\]

OpenStudy (anonymous):

Somewhat your final answer is triple of mine. What did you do? @dtan5457

OpenStudy (dtan5457):

\[\sqrt{\frac{ 1+\frac{ 4 }{ 5 } }{ 2 }}=\sqrt{\frac{ 9 }{ 10 }}=\frac{ 3\sqrt{10} }{ 10 }\]

OpenStudy (anonymous):

Why 1 + 4/5?

OpenStudy (dtan5457):

quadrant 1, positive, cos a=4/5 1+4/5

OpenStudy (dtan5457):

@Hoslos

OpenStudy (anonymous):

Ok. I am not sure about that identity, but more simply, you should first find ''a'' given in the formula cos a=4/5. Then plot this a in the formula of sin given by sin a/2.

OpenStudy (anonymous):

Or directly like this:\[\sin (\frac{ \cos^{-1} (\frac{ 4 }{ 5} )}{ 2 })\]

OpenStudy (dtan5457):

our answer is suppose to be the "exact answer" and have radicals, not be in complete decimal form